### v1

@@ 2.2 Gas radiative cooling and heating
Gas is allowed to cool down to $10\,\rm K$ through H and He collisions with a contribution from metals in the gas phase using rates tabulated by Sutherland & Dopita (1993) above $10^{4}\,\rm K$ and those from Dalgarno & McCray (1972) below $10^{4}\,\rm K$. Dust does not directly contribute to the gas cooling rates at low temperature, however, these rates can be significantly reduced in regions where the dust fraction over total metallicity is large, by decreasing the amount of metals in the gas phase available for gas cooling. The contribution of metals to the overall cooling rates is reduced by the corresponding amount of metals locked into the dust phase. We note that this approach is not consistent since the cooling curves, as tabulated in this version of ramses, are for a total metallicity with a solar composition of elements, while the amount of depletion is not uniform across all elements, and, hence, this should affect differently the cooling curves as a function of temperature. A more consistent approach will be explored using the prism ISM model for ramses from Katz et al. (2022) that computes cooling rates from the amount of each individual element and their ionisation state in the gas phase (Rodríguez Montero et al., in prep.).
At high temperatures, fast moving electrons frequently collide with grains, leading to an exchange of internal energy from the gas phase to the dust phase. Since dust radiates the stored energy into the IR, this leads to a net cooling for the gas phase. We use the corresponding cooling rates from Dwek & Werner (1981):
@@
where $x=2.71\times 10^{8}(a/\mu{\rm m})^{2/3}(T/{1\,\rm K})^{-1}$, and $x_{\rm max}=14000\left(a/\mu{\rm m}\right)^{2/3}$, which is introduced to impose a sharp cut-off to the dust cooling function at temperatures below $2\times 10^{4}\,\rm K$ in order to prevent the contribution to this cooling from dust in the cold phase. While the dust contribution to the gas cooling rate can largely exceed that of the fiducial gas cooling at $T>10^{6}\,\rm K$, thermal sputtering from ion collisions efficiently erodes dust grains (see Section 3.3), such that grains are destroyed on a timescale shorter than that of the gas cooling from dust (Guillard et al. 2009), which reduces the effective contribution to cooling at high temperatures (Montier & Giard 2004; Pointecouteau et al. 2009; Vogelsberger et al. 2019).
@@ 2.3 Star formation
Star formation follows a Schmidt law: $\dot{\rho}_{\star}=\varepsilon_{\star}{\rho_{\rm g}/t_{\rm ff}}$, where $\dot{\rho}_{\star}$ is the star formation rate mass density, $\rho_{\rm g}$ the gas mass density, $t_{\rm ff}$ the local free-fall time of the gas, and $\varepsilon_{\star}$ is the star formation efficiency. Star formation occurs in regions with gas number density above $n_{0}$ which varies with resolution ($n_{0}=0.6,2.5,10,40\,\rm H\,cm^{-3}$ for a minimum spatial resolution of respectively $\Delta x=72,36,18,9\,\rm pc$). This corresponds to a stellar mass resolution for newly formed stars of respectively $=7.4\times 10^{3},3.7\times 10^{3},1.9\times 10^{3},0.9\times 10^{3}\,\rm M_{\odot}$.
As in Dubois et al. (2021) (see for details of the implementation and the underlying parameters), the star formation efficiency varies with the virial parameter $\alpha_{\rm vir}=E_{\rm turb}/E_{\rm g}<1$, with $E_{\rm turb}$ and $E_{\rm g}$ the turbulent kinetic energy and the gravitational energy of the gas respectively, and with the turbulent Mach number $\mathcal{M}=\sigma_{\rm t}/c_{\rm s}$, where $\sigma_{\rm t}$ is the gas turbulent velocity and $c_{\rm s}$ is the sound speed. In practice, $\sigma_{\rm t}$ is computed by taking for each cell its own gas velocity and that of its neighboring cells into account, removing the local bulk (mass-weighted) mean velocity, and constructing $\sigma$ through $\sigma^{2}={\rm sum}(\nabla\otimes u{\rm d}x)^{2}$. In this theory (e.g. Krumholz & McKee 2005; Padoan & Nordlund 2011; Hennebelle & Chabrier 2011; Federrath & Klessen 2012), star formation (SF) is driven by how much gas mass passes a given density threshold. This amount is controlled by the level of turbulence and how much the molecular cloud is gravitationally bound. The unresolved density distribution can be described by a log-normal probability density function (PDF), with standard deviation given by $\mathcal{M}$ and $\alpha_{\rm vir}$: the more turbulent and bound the gas is, the larger the SF efficiency, reaching values as high as $\varepsilon_{\star}\sim 1$.
The mean SFR between 200 Myr and 400 Myr for the different simulated galaxies are $0.01\,\rm M_{\odot}\,yr^{-1}$, $0.03\,\rm M_{\odot}\,yr^{-1}$, $0.5\,\rm M_{\odot}\,yr^{-1}$, and $0.7\,\rm M_{\odot}\,yr^{-1}$ for respectively G8HG, G9LG, G9HG, and G10LG.
@@ 2.4 Stellar yields and feedback
We limit our simulations to individually track the amount of C, N, O, Mg, Si, Fe, S (which represents $\sim 85\%$ of the mass of solar heavy elements), of total metallicity Z, and H, but these can easily be extended to other individual elements. We assume an Chabrier (2005) initial mass function for the distribution of zero age star (ZAS) masses. Stellar gas plus dust yields for intermediate stars are those of Karakas (2010) for ZAS masses $M_{\rm ZAS}<8\,\rm M_{\odot}$ (asymptotic giant branch, AGB, stars), while yields from Limongi & Chieffi (2018) for massive stars $M_{\rm ZAS}\geq 8\,\rm M_{\odot}$ are employed. The yields for massive stars are parametrised by the rotation velocity of the star (either 0, 150 or $300\,\rm km\,s^{-1}$). We follow the initial distribution of rotational velocities as a function of the initial metallicity of the star from Prantzos et al. (2018) (obtained from multiple Milky Way chemical constraints), with the corresponding grid [^2] The range of available stellar yields for a given value of metallicity and rotational velocity is more limited but we simply bi-linearly interpolate the individual stellar yields from their data, which provides a finer sampled grid of IMF-weighted SSP yields as a function of metallicity. In addition, it should be noted that the yields of Limongi & Chieffi (2018) are given as a function of the Fe content in the ZAS. We prefer using the $Z$ metallicity content instead of the Fe content, which is more flexible. We have therefore converted their Fe content into $Z$ content by simply assuming solar ratios. of velocity-to-metallicity of ZAS: $V_{\rm ZAS}=150,100,50,50,50,50,50\,\rm km\,s^{-1}$ for $Z_{\rm ZAS}=10^{-3},10^{-2},10^{-1},10^{-0.6},10^{-0.3},1,10^{0.3}\,Z_{\odot}$ (where $Z_{\odot}=0.01345$ is the solar value of metallicity from Asplund et al. 2009). Since the Limongi & Chieffi (2018) stellar yields largely underproduce the amount of Mg, with respect to Si and solar abundances (Prantzos et al. 2018), we artificially boost up by a factor of 2 the amount of Mg released by massive stars. The mass distribution of stars is assumed to cover the $[0.1,100]\,\rm M_{\odot}$ mass range, assuming that massive stars only successfully explode in SNII and release their corresponding SN-processed yields for $M_{\rm ZAS}\leq 30\,\rm M_{\odot}$, while they also release mass through winds over their entire mass range ($[8,100]\,\rm M_{\odot}$). The death of each individual intermediate mass star follows the Padova stellar tracks (Girardi et al. 2000) with thermally pulsating AGB (Vassiliadis & Wood 1993). We further assume that the mass release of individual AGB stars happens instantaneously at the end of their evolution (pulsating evolution is not individually time-resolved).
@@
For type Ia SN (SNIa), we assume the delay time distribution of Maoz & Mannucci (2012) with a time-integrated number of SNIa of $N_{\rm Ia}(<t)=2.35\times 10^{-4}(\log(t/{\rm Gyr})+3)\,\rm M_{\odot}^{-1}$ (limiting the range of SNIa between 50 Myr and 13.7 Gyr), together with the stellar yields of Iwamoto et al. (1999) (using their W70 carbon-deflagration model).
The resulting mass release for each individual elements are shown in Appendix A.
@@
We simplify further the feedback from SNII by assuming that all mass and energy are released in one single explosive event after $5\,\rm Myr$ to maximise the impact of SNII feedback. We assume a Chabrier (2005) initial mass function, with a SNII explodability range of $M_{\rm ZAS}=[8,30]\rm M_{\odot}$ of canonical individual kinetic energy of $10^{51}\rm\,erg$, hence, corresponding to a SNII specific energy of $10^{49}\,\rm erg\,M_{\odot}^{-1}$. We include the SN feedback model from Kimm & Cen (2014), where the model follows either the energy conserving or momentum conserving phase of the explosion. If the SN explosion is still in the energy-conserving phase, only internal energy is given to the gas since the code will be able to handle the Sedov explosion and get the correct amount of momentum at the end of the Sedov phase. However, out of the energy-conserving phase, the correct amount of momentum is given to the gas. We also include a minimal model for UV radiation from OB young stars, by considering the larger amount of momentum SNII can impart to the gas thanks to the pre-heating by the UV-radiation (Geen et al. 2015).
@@ 3 Dust
The two bin size decomposition of Hirashita (2015) (see also the implementations of Aoyama et al. 2017, Gjergo et al. 2018, or Granato et al. 2021) is adopted for our subgrid model of dust evolution. These bins correspond to a small bin and a large bin of dust sizes (radii; $a$) of respectively $a_{\rm S}=5\,\rm nm$ and $a_{\rm L}=0.1\,\mu\rm m$. This decomposition allows to capture the bulk of the grain size distribution (Hirashita 2015), while having a negligible impact on the computational footprint (and on the memory load) of hydrodynamical simulations of galaxy formation.
@@
The equations of evolution of the dust mass content $D_{\rm i,j}$ of each grain size bin (with subscript $j=\rm S$ and L for respectively small and large grains) with chemical composition from the ‘key’ element $i$, which is C for carbonaceous grains, and Si for silicate grains is:
@@
In this work we neglect entirely the differential dynamical motion of dust with respect to gas dynamics, therefore, the dust content of each bin size and chemical composition can be treated as a passive variable that is transported along with the flow of gas in the exact same vein as for the transport of metals in the gas phase. This is possible due to the relatively short stopping time associated to Epstein aero-dynamical drag in dense gas $t_{\rm st}\simeq 0.5\,a_{0.1}(n/1\,{\rm cm^{-3}})^{-1}(c_{\rm s}/10\,\rm km\,s^{-1})^{-1}\,\rm Myr$ compared to dynamical time $t_{\rm dyn}\simeq 10\,(L/100\,\rm pc)(\sigma_{\rm gr}/10\,\rm km\,s^{-1})^{-1}\,\rm Myr$, where $\sigma_{\rm gr}$ is the grain velocity dispersion. This gives a diffusion length scale of $L\simeq 5\,a_{0.1}(n/1\,{\rm cm^{-3}})^{-1}(\sigma_{\rm gr}/c_{\rm s})\,\rm pc$, which is in general below the resolution scale of this work. We defer the investigation of the diffusion of dust grains on cloud scales to future work.
@@ 3.1 Stellar production
Similarly to silicates, the amount of carbonaceous dust, and the condensation efficiency $\delta_{\rm C}$ vary with the nature of the stellar ejecta. It is assumed to be $\delta_{\rm C}^{\rm SNII}=0.5$ and $\delta_{\rm C}^{\rm SNIa}=0.5$, while for AGB stars this efficiency depends on the ratio of C/O in the ejecta. If the ratio of C/O $<1$, then all the C elements forms into CO molecules that are not available for dust $\delta_{\rm C}^{\rm AGB,C/O<1}=0$, while for C/O $\geq 1$ the efficiency is 1 with a mass released into C dust:
@@
The production of dust in stellar ejecta is balanced between a population of small and large grains, i.e. $\dot{D}_{{\rm ej},i,{\rm S}}=f_{{\rm ej},i,{\rm S}}\dot{D}_{{\rm ej},i}$ and $\dot{D}_{{\rm ej},i,{\rm L}}^{k}=(1-f_{{\rm ej},i,{\rm S}})\dot{D}_{{\rm ej},i}$, where $f_{{\rm ej},i,{\rm S}}$ is the fraction of small grains released for a given chemical type $i$ (carbonaceous or silicate grains). For sake of simplicity, we assume that all the dust released in the ejecta is condensed into the large grain size population only, thus, $f_{{\rm ej},i,{\rm S}}=0$ (see appendix B for the predictions of extinction curves with different values of $f_{{\rm ej},i,{\rm S}}$, and the discussion in section 6).
Finally, it is important to note that the predicted and observed dust condensation efficiencies are very uncertain and can vary by an order of magnitude (see Schneider & Maiolino 2023).
3.2 SN-driven destruction

### v2

@@ 2.2 Gas radiative cooling and heating
Gas was allowed to cool down to $10\,\rm K$ through H and He collisions with a contribution from metals in the gas phase using rates tabulated by Sutherland & Dopita (1993) above $10^{4}\,\rm K$ and those from Dalgarno & McCray (1972) below $10^{4}\,\rm K$. Dust did not directly contribute to the gas cooling rates at low temperature, however, these rates can be significantly reduced in regions where the dust fraction over total metallicity is large, by decreasing the amount of metals in the gas phase available for gas cooling. The contribution of metals to the overall cooling rates was reduced by the corresponding amount of metals locked into the dust phase. We note that this approach is not consistent since the cooling curves, as tabulated in this version of ramses, are for a total metallicity with a solar composition of elements, while the amount of depletion is not uniform across all elements, and, hence, this should affect differently the cooling curves as a function of temperature. A more consistent approach will be explored using the prism ISM model for ramses from Katz et al. (2022) that computes cooling rates from the amount of each individual element and their ionisation state in the gas phase (Rodríguez Montero et al., in prep.).
At high temperatures, fast moving electrons frequently collide with grains, leading to an exchange of internal energy from the gas phase to the dust phase. Since dust radiates the stored energy into the IR, this leads to a net cooling for the gas phase. We used the corresponding cooling rates from Dwek & Werner (1981):
@@
where $x=2.71\times 10^{8}(a/\mu{\rm m})^{2/3}(T/{1\,\rm K})^{-1}$, and $x_{\rm max}=14000\left(a/\mu{\rm m}\right)^{2/3}$, which was introduced to impose a sharp cut-off to the dust cooling function at temperatures below $2\times 10^{4}\,\rm K$ in order to prevent the contribution to this cooling from dust in the cold phase. While the dust contribution to the gas cooling rate can largely exceed that of the fiducial gas cooling at $T>10^{6}\,\rm K$, thermal sputtering from ion collisions efficiently erodes dust grains (see Section 3.3), such that grains are destroyed on a timescale shorter than that of the gas cooling from dust (Guillard et al. 2009), which reduces the effective contribution to cooling at high temperatures (Montier & Giard 2004; Pointecouteau et al. 2009; Vogelsberger et al. 2019).
@@ 2.3 Star formation
Star formation followed a Schmidt law: $\dot{\rho}_{\star}=\varepsilon_{\star}{\rho_{\rm g}/t_{\rm ff}}$, where $\dot{\rho}_{\star}$ is the star formation rate mass density, $\rho_{\rm g}$ the gas mass density, $t_{\rm ff}$ the local free-fall time of the gas, and $\varepsilon_{\star}$ is the star formation efficiency. Star formation occurred in regions with gas number density above $n_{0}$ which varied with resolution ($n_{0}=0.6,2.5,10,40\,\rm H\,cm^{-3}$ for a minimum spatial resolution of respectively $\Delta x=72,36,18,9\,\rm pc$). This corresponds to a stellar mass resolution for newly formed stars of respectively $=7.4\times 10^{3},3.7\times 10^{3},1.9\times 10^{3},0.9\times 10^{3}\,\rm M_{\odot}$.
As in Dubois et al. (2021) (see for details of the implementation and the underlying parameters), the star formation efficiency varied with the virial parameter $\alpha_{\rm vir}=2E_{\rm turb}/|E_{\rm g}|<1$, with $E_{\rm turb}$ and $E_{\rm g}$ the turbulent kinetic energy and the gravitational energy of the gas respectively, and with the turbulent Mach number $\mathcal{M}=\sigma_{\rm t}/c_{\rm s}$, where $\sigma_{\rm t}$ is the gas turbulent velocity and $c_{\rm s}$ is the sound speed. In practice, $\sigma_{\rm t}$ was computed by taking for each cell its own gas velocity and that of its neighbouring cells into account, removing the local bulk (mass-weighted) mean velocity, and constructing $\sigma$ through $\sigma^{2}={\rm sum}(\nabla\otimes u{\rm d}x)^{2}$. In this theory (e.g. Krumholz & McKee 2005; Padoan & Nordlund 2011; Hennebelle & Chabrier 2011; Federrath & Klessen 2012), star formation (SF) is driven by how much gas mass passes a given density threshold. This amount is controlled by the level of turbulence and how much the molecular cloud is gravitationally bound. The unresolved density distribution can be described by a log-normal probability density function (PDF), with standard deviation given by $\mathcal{M}$ and $\alpha_{\rm vir}$: the more turbulent and bound the gas is, the larger the SF efficiency, reaching values as high as $\varepsilon_{\star}\sim 1$. The mean SFR between 200 Myr and 400 Myr for the different simulated galaxies are $0.01\,\rm M_{\odot}\,yr^{-1}$, $0.03\,\rm M_{\odot}\,yr^{-1}$, $0.5\,\rm M_{\odot}\,yr^{-1}$, and $0.7\,\rm M_{\odot}\,yr^{-1}$ for respectively G8HG, G9LG, G9HG, and G10LG.
@@ 2.4 Stellar yields and feedback
We limited our simulations to individually track the amount of C, N, O, Mg, Si, Fe, S (which represented $\sim 85\%$ of the mass of solar heavy elements), of total metallicity Z, and H, but these can easily be extended to other individual elements. We assumed a Chabrier (2005) initial mass function for the distribution of zero age star (ZAS) masses. Stellar gas plus dust yields for intermediate stars were those of Karakas (2010) for ZAS masses $M_{\rm ZAS}<8\,\rm M_{\odot}$ (asymptotic giant branch, AGB, stars), while yields from Limongi & Chieffi (2018) for massive stars $M_{\rm ZAS}\geq 8\,\rm M_{\odot}$ were employed. The yields for massive stars were parametrised by the rotation velocity of the star (either 0, 150 or $300\,\rm km\,s^{-1}$). We followed the initial distribution of rotational velocities as a function of the initial metallicity of the star from Prantzos et al. (2018) (obtained from multiple Milky Way chemical constraints), with the corresponding grid [^2] The range of available stellar yields for a given value of metallicity and rotational velocity is more limited but we simply bi-linearly interpolated the individual stellar yields from their data, which provided a finer sampled grid of IMF-weighted SSP yields as a function of metallicity. In addition, it should be noted that the yields of Limongi & Chieffi (2018) are given as a function of the Fe content in the ZAS. We prefered using the $Z$ metallicity content instead of the Fe content, which is more flexible. We therefore converted their Fe content into $Z$ content by simply assuming solar ratios. of velocity-to-metallicity of ZAS: $V_{\rm ZAS}=150,100,50,50,50,50,50\,\rm km\,s^{-1}$ for $Z_{\rm ZAS}=10^{-3},10^{-2},10^{-1},10^{-0.6},10^{-0.3},1,10^{0.3}\,Z_{\odot}$ (where $Z_{\odot}=0.01345$ is the solar value of metallicity from Asplund et al. 2009). Since the Limongi & Chieffi (2018) stellar yields largely underproduce the amount of Mg, with respect to Si and solar abundances (Prantzos et al. 2018), we artificially boosted up by a factor of 2 the amount of Mg released by massive stars. The mass distribution of stars was assumed to cover the $[0.1,100]\,\rm M_{\odot}$ mass range, assuming that massive stars only successfully explode in SNII and release their corresponding SN-processed yields for $M_{\rm ZAS}\leq 30\,\rm M_{\odot}$, while they also release mass through winds over their entire mass range ($[8,100]\,\rm M_{\odot}$). The death of each individual intermediate mass star followed the Padova stellar tracks (Girardi et al. 2000) with thermally pulsating AGB (Vassiliadis & Wood 1993). We further assumed that the mass release of individual AGB stars happens instantaneously at the end of their evolution (pulsating evolution is not individually time-resolved).
@@
For type Ia SN (SNIa), we assumed the delay time distribution of Maoz & Mannucci (2012) with a time-integrated number of SNIa of $N_{\rm Ia}(<t)=2.35\times 10^{-4}(\log(t/{\rm Gyr})+3)\,\rm M_{\odot}^{-1}$ (limiting the range of SNIa between 50 Myr and 13.7 Gyr), together with the stellar yields of Iwamoto et al. (1999) (using their W70 carbon-deflagration model). The resulting mass release for each individual elements are shown in Appendix A.
@@
We simplified further the feedback from SNII by assuming that all mass and energy are released in one single explosive event after $5\,\rm Myr$ to maximise the impact of SNII feedback. We assumed a Chabrier (2005) initial mass function, with a SNII explodability range of $M_{\rm ZAS}=[8,30]\rm M_{\odot}$ of canonical individual kinetic energy of $10^{51}\rm\,erg$, hence, corresponding to a SNII specific energy of $10^{49}\,\rm erg\,M_{\odot}^{-1}$. We included the SN feedback model from Kimm & Cen (2014), where the model follows either the energy conserving or momentum conserving phase of the explosion. If the SN explosion is still in the energy-conserving phase, only internal energy is given to the gas since the code will be able to handle the Sedov explosion and get the correct amount of momentum at the end of the Sedov phase. However, out of the energy-conserving phase, the correct amount of momentum is given to the gas. We also included a minimal model for UV radiation from OB young stars, by considering the larger amount of momentum SNII can impart to the gas thanks to the pre-heating by the UV-radiation (Geen et al. 2015).
@@ 3 Dust
The two bin size decomposition of Hirashita (2015) (see also the implementations of Aoyama et al. 2017, Gjergo et al. 2018, or Granato et al. 2021) was adopted for our subgrid model of dust evolution. These bins correspond to a small bin and a large bin of dust sizes (radii; $a$) of respectively $a_{\rm S}=5\,\rm nm$ and $a_{\rm L}=0.1\,\mu\rm m$. This decomposition allows to capture the bulk of the grain size distribution (Hirashita 2015), while having a negligible impact on the computational footprint (and on the memory load) of hydrodynamical simulations of galaxy formation.
@@
The equations of evolution of the dust mass content $D_{\rm i,j}$ of each grain size bin (with subscript $j=\rm S$ and L for respectively small and large grains) with chemical composition from the ‘key’ element $i$, which is C for carbonaceous grains, and Si for silicate grains, are:
@@
In this work we neglected entirely the differential dynamical motion of dust with respect to gas dynamics, therefore, the dust content of each bin size and chemical composition can be treated as a passive variable that is transported along with the flow of gas in the exact same vein as for the transport of metals in the gas phase. This is possible due to the relatively short stopping time associated to Epstein aero-dynamical drag in dense gas $t_{\rm st}\simeq 0.5\,a_{0.1}(n/1\,{\rm cm^{-3}})^{-1}(c_{\rm s}/10\,\rm km\,s^{-1})^{-1}\,\rm Myr$ compared to dynamical time $t_{\rm dyn}\simeq 10\,(L/100\,\rm pc)(\sigma_{\rm gr}/10\,\rm km\,s^{-1})^{-1}\,\rm Myr$, where $\sigma_{\rm gr}$ is the grain velocity dispersion. This gives a diffusion length scale of $L\simeq 5\,a_{0.1}(n/1\,{\rm cm^{-3}})^{-1}(\sigma_{\rm gr}/c_{\rm s})\,\rm pc$, which is in general below the resolution scale of this work. We defer the investigation of the diffusion of dust grains on cloud scales to future work.
@@ 3.1 Stellar production
Similarly to silicates, the amount of carbonaceous dust and the condensation efficiency $\delta_{\rm C}$ varied with the nature of the stellar ejecta. It was assumed to be $\delta_{\rm C}^{\rm SNII}=0.5$ and $\delta_{\rm C}^{\rm SNIa}=0.5$, while for AGB stars this efficiency depended on the ratio of C/O in the ejecta. If the ratio of C/O $<1$, then all the C elements formed into CO molecules that were not available for dust $\delta_{\rm C}^{\rm AGB,C/O<1}=0$, while for C/O $\geq 1$ the efficiency was 1 with a mass released into C dust:
@@
The production of dust in stellar ejecta is balanced between a population of small and large grains, i.e. $\dot{D}_{{\rm ej},i,{\rm S}}=f_{{\rm ej},i,{\rm S}}\dot{D}_{{\rm ej},i}$ and $\dot{D}_{{\rm ej},i,{\rm L}}^{k}=(1-f_{{\rm ej},i,{\rm S}})\dot{D}_{{\rm ej},i}$, where $f_{{\rm ej},i,{\rm S}}$ is the fraction of small grains released for a given chemical type $i$ (carbonaceous or silicate grains). For sake of simplicity, we assumed that all the dust released in the ejecta is condensed into the large grain size population only, thus, $f_{{\rm ej},i,{\rm S}}=0$ (see appendix B for the predictions of extinction curves with different values of $f_{{\rm ej},i,{\rm S}}$, and the discussion in section 6). Finally, it is important to note that the predicted and observed dust condensation efficiencies are very uncertain and can vary by an order of magnitude (see Schneider & Maiolino 2023).
3.2 Supernova-driven destruction